OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = 5/7x - 12 A. -5x + 7y = -12 B. -5x - 7y = -84 C. -5x + 7y = -84 D. 5x + 7y = -84

OpenStudy (anonymous):

first Get x and y on the same side so that your equation has this form ax+by=c

OpenStudy (anonymous):

Follow the procedure I described on your previous question.

OpenStudy (anonymous):

okay , one second

OpenStudy (anonymous):

so it would look like 5/7x + y = -12 ?

OpenStudy (anonymous):

Since its positive on the right side, you subtract it to move it to the left -5/7x+y=-12 now multiply everything by the lcd which is 7

OpenStudy (anonymous):

but the 12 is negative?

OpenStudy (anonymous):

Yes it is. What do you mean?

OpenStudy (anonymous):

you said since its positive on the right side... sorry ,what did you mean by that then?

OpenStudy (anonymous):

oh i meant the positive 5/7x, you subtract it from both sides since its positive on the right hand side

OpenStudy (anonymous):

how do i do that? sorry , i'm confused.

OpenStudy (anonymous):

its like when you're solving an equation, when something is being added to one side, to move it to the other side you subtract it from both sides right? so with y=5/7x-12 5/7x is positive on the right side and you want it on the left so you subtract it -5/7x+y=-12

OpenStudy (anonymous):

ohh , okay i understand, now multiply -5/7x times 7 and -12 times 7?

OpenStudy (anonymous):

Multiply every term by 7

OpenStudy (anonymous):

-5x+7y=-84?

OpenStudy (anonymous):

The first coefficient may not be negative.

OpenStudy (anonymous):

when i multiplied -5/7x times 7 it gave me -5 unless i put it in incorrectly.

OpenStudy (anonymous):

I forgot the first coefficient couldnt be negative =p so instead of 7, you would multiply everything by -7

OpenStudy (anonymous):

5o then D?

OpenStudy (anonymous):

?? if i do that the answers dont fit my answer choices.....

OpenStudy (anonymous):

the answer was C.